OpenStudy (trojanpoem):

How can I put put this equation x^3+x^2-x+15 = 0 One of this equation answers is -3 So i wanna turn it into something like this (x+3)*Z X is the rest of the equation. If it's not clear , just tell me.

OpenStudy (anonymous):

Dude, you can just Google that. You'll find answers how to do it.

OpenStudy (trojanpoem):

What to google for ?

OpenStudy (michele_laino):

Applying the Ruffini rule you get: \[x ^{3}-x ^{2}+x+15=(x+3)*(x ^{2}-4x+5)\]

OpenStudy (michele_laino):

Sorry I have made an error!

OpenStudy (trojanpoem):

You meant -2x right ?

OpenStudy (trojanpoem):

Then what I am looking for is Ruffini rule :D

OpenStudy (michele_laino):

Applying the Ruffini rule, I get: \[x ^{3}+x ^{2}-x+15=(x+3)*(x ^{2}-2x+5)\]